# FedCLEAN: byzantine defense by CLustering Errors of Activation maps in Non-IID federated learning environments

### v1

@@ Abstract
Federated Learning (FL) enables clients to collaboratively train a global model using their local datasets while reinforcing data privacy. However, FL is susceptible to poisoning attacks. Existing defense mechanisms assume that clients’ data are independent and identically distributed (IID), making them ineffective in real-world applications where data are non-IID. This paper presents FedCLEAN, the first defense capable of filtering attackers’ model updates in a non-IID FL environment. The originality of FedCLEAN is twofold. First, it relies on a client confidence score derived from the reconstruction errors of each client’s model activation maps for a given trigger set, with reconstruction errors obtained by means of a Conditional Variational Autoencoder trained according to a novel server-side strategy. Second, we propose an ad-hoc trust propagation algorithm based on client scores, which allows building a cluster of benign clients while flagging potential attackers. Experimental results on the datasets MNIST and FashionMNIST demonstrate the robustness of FedCLEAN against Byzantine attackers in non-IID scenarios and a close-to-zero benign client misclassification rate, even in the absence of an attack.
Keywords:
Poisoning attacks Byzantine attacks Non-IID Federated Learning Conditional variational autoencoders Clustering
@@ 1 Introduction
In recent years, Federated Learning (FL) has attracted considerable interest as a privacy-preserving machine learning paradigm [40, 32, 25] in several domains, such as healthcare [36], transportation [36], and personal devices [28]. In a typical FL scenario, several clients collectively contribute to training a global model. At first, a central server initiates the global model and sends it to the clients so that they can train it locally on their private data. Model updates are sent back to the server, which aggregates them into a new global model. These two steps are repeated until the global model converges. As clients’ data remain stored locally, privacy is ensured [42]. However, even though FL systems preserve privacy, their decentralized nature makes them vulnerable to poisoning attacks, where one or several malicious clients try to corrupt or prevent the training session by sending malicious updates [50]. Some defenses try to mitigate these attacks using secure aggregation algorithms with the idea of reducing the impact of bad updates [41, 8, 26]. Nevertheless, these solutions remain limited, being, for instance, not resistant to attacks that can destroy the global model after just one round of a successful attack, like Byzantine attacks [12]. This has led to the proposal of FL defense systems devoted to the detection of malicious updates before the aggregation stage. Solutions proposed to date [24, 11, 10, 46, 6] have been proven to be Byzantine attack-robust under the hypothesis of Independent and Identically Distributed (IID) data, i.e., when training data is homogeneously distributed among clients in terms of the number of samples and class balance. However, such systems do not perform well in non-IID contexts [2], which are more likely encountered in real-life scenarios. In the following, we introduce these solutions and their limitations, then present our contribution: FedCLEAN, the first FL defense system adapted to non-IID contexts.
@@ 1.1 Related Works
As stated above, the first defenses against malicious updates were the development of robust aggregation algorithms. FedAvg [41] is one of the earliest and most commonly used aggregation methods in FL. It consists of using a weighted average of the updates sent by clients to actualize the global model. Clients’ dataset sizes are used as weights. Doing so dilutes the effect of sparse attackers but is not sufficient when their number increases. It is also not robust to FL attacks that amplify updates to dominate the global average [48], and ineffective in non-IID scenarios where an attacker has a larger dataset compared to benign clients. GeoMed [26] uses the geometric median of clients’ updates as a global update. KRUM [8] uses one of the local updates to generate a global model update. The chosen local update minimizes the sum of distances to its nearest neighbors. Other aggregation schemes have been proposed since [38, 34, 1, 23], but they focus on training performance rather than security. All of these methods have been shown to be ineffective under Byzantine attacks [24].
@@
To better respond to attacks, "detect and remove" defense systems have been recently proposed. They aim at excluding malicious updates before the aggregation step. One first class of methods regroups systems which rely on anomaly detection techniques based on autoencoders. Li et al. [37] were the first to propose such a defense. It trains a copy of the global model on the test dataset on the server side, while training at the same time a variational autoencoder (VAE) on surrogate vectors randomly sampled from the weights of this model at each epoch. Although this solution performs better than aggregation-based defenses under Byzantine attacks, it requires training a VAE on an entire server-side dataset to achieve good results. This is quite unrealistic in real-life scenarios, as the server most likely won’t have access to a significant training dataset. FedCVAE [24] later proposed an improvement using an untrained conditional variational autoencoder (CVAE). When the central server receives updates from the clients, the defense system computes the geometric median (GM) of surrogate vectors from each update. Then it normalizes these vectors by subtracting their GM value. These normalized vectors are then inputted into the CVAE, conditioned on the current round number in the FL training. As the CVAE is not pre-trained, only inputs near zero yield reconstruction errors close to zero, and vice versa. FedCVAE shows some defense results in a non-IID scenario but only against attacks exaggerated in magnitude and thus fairly easy to detect. More recently, FedCAM [6], a CVAE anomaly detection-based approach, achieves the latest state-of-the-art results against Byzantine attacks in IID scenarios. Rather than using model weights, it works with their activation maps for a given trigger set on which a CVAE is trained with GeoMed as a regularization. However, FedCAM only works in IID scenarios, as we will see in Section 5.
@@
A last class corresponds to clustering-based approaches. They have shown promising results on anomaly detection without needing autoencoders. Among recent ones, we can cite DeepSight [46]. It uses a model inspection technique and a clustering scheme to create clusters of updates, and to pick out those with the best scores. DeepSight successfully stops backdoor attackers [27], even in the non-IID case, by selecting groups with minimal attacker influence. However, it does not work for Byzantine attacks where a single successful breach is sufficient to make the global model convergence fail. More recently, FLEDGE [10] used a combination of cosine similarity and clustering to filter out attacks. It provides results against Byzantine attacks but only a fraction of their data is non-IID.
@@ 1.2 Contribution
In this work, we propose FedCLEAN, an extension of FedCAM, a solution introduced by Bellafqira et al. [6]. Similar to FedCAM, FedCLEAN analyzes the GeoMed normalized CVAE reconstruction errors of clients’ activation maps for given inputs of a trigger set. Clients with high reconstruction errors are rejected based on an anomaly detection strategy. While FedCAM achieves top byzantine robustness compared to concurrent methods in IID scenarios [6], it fails to achieve good results in the context of non-IID data, as we will see in Section 5. We identified two main issues causing this failure, which we address with FedCLEAN. First, the CVAE training method proposed in FedCAM has poor generalization in non-IID contexts, as client updates are more heterogeneous. Specifically, clients with the most unique data often fall out of distribution for CVAE and end up with similar error scores to attackers. To overcome this issue, FedCLEAN employs a different CVAE training process that accounts for the temporal evolution of FL training with warmup epochs and better captures the statistical properties of benign updates using an annealing technique and layer selection. This ensures that all benign client activation maps, even outliers, are correctly reconstructed. Second, FedCAM relies on the mean of reconstruction errors as a dynamic error threshold. This approach risks allowing the selection of some attackers’ updates for the aggregation step if the average is high enough. It is important to note that retaining even one Byzantine attacker in a round is sufficient to destroy the global model. Moreover, this strategy also impacts the convergence or accuracy of the global model in the absence of attackers, as it eliminates around half of the participants in each round. In the non-IID case, clients are less redundant compared to IID data. To address this problem, FedCLEAN introduces a clustering algorithm that iteratively aggregates and concentrates benign candidates into a trusted cluster, enabling a more refined selection process. The rest of this paper is organized as follows. Section 2 provides an overview of FL technical background, FL attacks, and useful components of our solution. Section 3 presents the federated context we considered and its threat model. Section 4 introduces our proposal, FedCLEAN, followed by Section 5, which details our benchmark and provides an analysis of the experimental results. Finally, we conclude the paper in Section 6.
@@ 2.1.1 A Common FL Scenario
In this work, we considered an FL setting where $N$ clients $\{C_{k}\}_{k=1..N}$ with their datasets $\{D_{k}\}_{k=1..N}$ participate in a training session orchestrated by a central server $S$ to train a global model $M_{G}$. During each FL round $t$, $S$ scatters the current global model $M_{G}^{t}$ to a subset of $U$ randomly selected clients. Each client updates the received model on its local dataset for a number of local epochs and sends back the updated model to $S$. The server then uses an aggregation algorithm such as FedAvg [41] to get an updated version of the global model $M_{G}^{t+1}$ from the clients’ updated models $\{M_{u}^{t}\}_{u=1..U}$. The above training steps are then repeated until convergence or a stopping criterion is met.
@@ 2.1.2 Non-IID Data in Federated Learning
In practice, feature and quality skew are intrinsic properties of data and are hard to simulate for an experiment if the dataset does not provide a natural partition. So, most works that address Non-IID FL use a combination of label and quantity skew by choosing a distribution of data over clients that ensures heterogeneity.
2.1.3 Attacks in Federated Learning
Federated Learning is known to be vulnerable to several attacks [33]. During training, these attacks involve an adversary that compromises one or several clients and tries to use them to disrupt the global model. Herein, two main approaches can be distinguished: "model poisoning", where the attacker can directly manipulate the compromised clients’ models and send malicious updates to the server; and "data poisoning", where the attacker indirectly manipulates learning by corrupting the clients’ local datasets. Based on the attacker’s objective, these attacks can further be divided into two sub-classes:
  • Untargeted poisoning attacks - where the attacker’s goal is to drop the global model accuracy on any future inputs by preventing its convergence. It is generally very hard for a machine learning model to recover from a bad convergence, and one attack in an unprotected FL round is often sufficient for an untargeted poisoning attack to succeed. Such attacks are also referred to as "Byzantine attacks".
  • Targeted poisoning attacks - also called "Backdoor attacks", where the attacker aims at preserving the accuracy of the model on its main task while simultaneously training it on a second malicious task, introducing a backdoor. The most common form of backdoor attacks involves inserting a trigger in the training dataset and training the model to output a specific response when the trigger data are given as an input.
It is important to notice that making a successful backdoor attack is very hard in Non-IID scenarios [27]. Controlling just some of the FL clients with non-IID data, the attacker generally has access only to a subpart of the training data, which in some cases is not sufficiently representative for injecting a robust backdoor to benign clients’ updates [39]. Moreover, the backdoor updates are easily overwritten by benign clients’ ones, as shown in [39]. This is the reason why in this work, we decided to focus on defending against Byzantine attacks, which currently pose a greater threat in FL.
@@ 2.2 Conditional Variational Autoencoders (CVAE)
Autoencoders (AEs) [30] are a class of unsupervised generative neural networks that aim to reconstruct a given input from a latent representation encoding. AEs are made of: i) an encoder $f$ that maps input vectors from the initial feature space to a lower-dimensionality representation space called the latent space; and ii) a decoder that retrieves input vectors from their latent representations. The objective of an autoencoder is to minimize the reconstruction error (RE) between the initial and reconstructed vectors. They have been historically used for dimensionality reduction [4, 30], model pre-training [20, 7], and generative tasks [52]. An important property of AEs is that, when applied to out-of-distribution vectors that deviate from the original training distribution, these deviations are amplified in the latent space, and much higher reconstruction errors are observed. This makes AEs considered as a valuable tool for anomaly detection in a wide range of applications [43].
Autoencoders having a fixed-dimensional latent space show some limitations for datasets with varying complexities. This has motivated the introduction of Variational Autoencoders (VAEs) [35] with a variational latent space based on Bayesian inference methods [18]. Conditional-VAEs (CVAEs) [45] are an extension of VAEs that leverages conditional probabilities to guide the encoding and decoding process following a provided condition. In practice, what differentiates a CVAE from a traditional VAE is the addition of a condition vector $y$ as an input of both the encoder and decoder (see Fig. 1). Thus, CVAEs enhance the capabilities of VAEs by making them compatible with supervised learning and allowing them to work in a task-guided manner [16]. They have been extensively used for a number of applications, including anomaly detection and detecting attackers in federated learning [24, 6, 11].
@@ 2.3 GeoMed and Activation Maps
In a typical deep learning model, the data is processed through feed-forward layers. Each hidden layer takes in the previous layer’s output, applies a linear transformation parametrized by a set of weights (e.g., a convolution or weighted sum), and then applies a nonlinearity function. The output of the layer’s weighted transformation is referred to as the activations of that layer. "Activation maps", or the set of activation maps extracted from a model for a given input, are characteristic of a model and the way it processes data. They have been used for applications such as explainability [47] and model watermarking [15, 5, 56]. It has been observed that in many real-life applications, the activation maps of a model tend to follow an intrinsic distribution [15, 44].

### v2

# FL-CLEANER: byzantine and backdoor defense by CLustering Errors of Activation maps in Non-iid fedErated leaRning

@@ Abstract
Federated Learning (FL) enables clients to collaboratively train a global model using their local datasets while reinforcing data privacy, but it is prone to poisoning attacks. Existing defense mechanisms assume that clients’ data are independent and identically distributed (IID), making them ineffective in real-world applications where data are non-IID. This paper presents FL-CLEANER, the first defense capable of filtering both byzantine and backdoor attackers’ model updates in a non-IID FL environment. The originality of FL-CLEANER is twofold. First, it relies on a client confidence score derived from the reconstruction errors of each client’s model activation maps for a given trigger set, with reconstruction errors obtained by means of a Conditional Variational Autoencoder trained according to a novel server-side strategy. Second, it uses an original ad-hoc trust propagation algorithm we propose. Based on previous client scores, it allows building a cluster of benign clients while flagging potential attackers. Experimental results on the datasets MNIST and FashionMNIST demonstrate the efficiency of FL-CLEANER against Byzantine attackers as well as to some state-of-the-art backdoors in non-IID scenarios; it achieves a close-to-zero (<1%) benign client misclassification rate, even in the absence of an attack, and achieves strong performance compared to state of the art defenses.
Keywords Poisoning attacks, Byzantine attacks, Backdoor attacks, Non-IID Federated Learning, Conditional variational autoencoders, Clustering.
@@ 1 Introduction
In recent years, Federated Learning (FL) has attracted considerable interest as a privacy-preserving machine learning paradigm [1, 2, 3] in several domains, such as healthcare [4], transportation [4], and personal devices [5]. In a typical FL scenario, several clients collectively contribute to training a global model. At first, a central server initiates the global model and sends it to the clients so that they can train it locally on their private data. Model updates, in the form of weights, are sent back to the server which aggregates them into a new global model. These two steps are repeated until the global model converges. As clients’ data remain stored locally, privacy is ensured [6]. However, even though FL systems preserve privacy, their decentralized nature makes them vulnerable to poisoning attacks, where one or several malicious clients try to corrupt or prevent the training session by sending malicious updates [7]. Some defenses try to mitigate the impact of these attacks using secure aggregation algorithms with the idea of reducing the impact of bad updates [8, 9, 10]. Nevertheless, these solutions remain limited. They are, for instance, not resistant to attacks able to destroy the global model after just one successful FL round of attack, known as Byzantine attacks [11]. This has led to the proposal of FL defense systems devoted to the detection of malicious updates before the aggregation stage. Solutions proposed to date [12, 13, 14, 15, 16] have been proven to be Byzantine attack-robust under the hypothesis of Independent and Identically Distributed (IID) data, i.e., when training data is homogeneously distributed among clients in terms of the number of samples and class balance. However, such systems do not perform well in non-IID contexts [17], which are more likely encountered in real-life scenarios. In the following, we introduce these solutions pointing out their limitations before then presenting our contribution: FL-CLEANER, the first FL defense system adapted to non-IID contexts. Before concluding, we will also experimentally show that FL-CLEANER: i) effectively blocks all byzantine attacks in a non-IID setting, all while having near zero false positives; ii) mitigates the effects of backdoor attacks.
Figure 1: A common online FL scenario with a central server
@@ 1.1 Related Works
As stated above, the first defenses against malicious clients’ updates were the development of robust aggregation algorithms. FedAvg [8] is one of the earliest and most commonly used aggregation methods in FL. It consists of using a weighted average of the updates sent by clients to actualize the global model. Clients’ dataset sizes are used as weights. Doing so dilutes the effect of attackers when they are few in number but is not sufficient when their number increases. It is also not robust to FL attacks that amplify updates to dominate the global average [18], and ineffective in non-IID scenarios especially when an attacker has a larger dataset compared to benign clients. GeoMed [10] uses the geometric median of clients’ updates as a global update. KRUM [9] uses one of the local updates to generate a global model update. The chosen local update minimizes the sum of distances to its nearest neighbors. Other aggregation schemes have been proposed since [19, 20, 21, 22], but they focus on training performance rather than security. All of these methods have been shown to be ineffective under Byzantine attacks [12].
@@
To better respond to attacks, "detect and remove" defense systems have been recently proposed. They aim at excluding malicious updates before the aggregation step. One first class of methods regroups systems which rely on anomaly detection techniques based on autoencoders. Li et al. [23] were the first to propose such a defense. It trains a copy of the global model on the test dataset on the server side, while training at the same time a variational autoencoder (VAE) on surrogate vectors (randomly sampled weights from the model) at each epoch. Although this solution performs better than aggregation-based defenses under Byzantine attacks, it requires training a VAE on an entire server-side dataset to achieve good results. This is quite unrealistic in real-life scenarios, as the server most likely won’t have access to a significant training dataset. FedCVAE [12] later proposed an improvement using an untrained conditional variational autoencoder (CVAE). When the central server receives updates from the clients, the defense system computes the geometric median (GM) of surrogate vectors from each update. Then it normalizes these vectors by subtracting their GM value. These normalized vectors are then inputted into the CVAE, conditioned on the current round number in the FL training. As the CVAE is not pre-trained, only inputs near zero yield reconstruction errors close to zero, and vice versa. FedCVAE shows some defense results in a non-IID scenario but only against attacks exaggerated in magnitude and thus fairly easy to detect. They also provide results on simple backdoor attacks. More recently, FedCAM [16], a CVAE anomaly detection-based approach, achieves the latest state-of-the-art results against Byzantine attacks in IID scenarios. Rather than using model weights, it works with their activation maps for a given trigger set on which a CVAE is trained with GeoMed as a regularization. However, FedCAM only works in IID scenarios, as we will see in Section 5.
@@
A last class corresponds to clustering-based approaches. They have shown promising results on anomaly detection without needing autoencoders. Among recent ones, we can cite FLEDGE [14] which used a combination of cosine similarity and Kernel Density Estimation to filter out attacks. It focuses on backdoor attacks but provides results against Byzantine attacks but only a fraction of their data is non-IID. Finally, Sharma et al. [25] recently proposed a defense based on graph representation combined with Graph Attention Networks. Beyond the fact they require pre-training a model on both attacker and benign client examples, its authors provide only results for one byzantine attack, being sign-flipping attack.
In addition to these defenses, several others exist that specifically target backdoor attacks [26]. Defenses such as FoolsGold [27], DeepSight [15], FLAME [28] and FLDetector [29]rely on filtering out attackers by finding clients that present similar behavior. Other propositions try to remove the effect of the backdoor after the aggregation step, such as [30] and [31]. However, these propositions have not been tested or proven effective against byzantine attacks.
@@ 1.2 Contribution
In this work, we propose FL-CLEANER, an extension of FedCAM, introduced in [16], which was only devoted to byzantine attack detection in IID settings. Similarly to FedCAM, FL-CLEANER analyzes the GeoMed normalized CVAE reconstruction errors of clients’ activation maps for given inputs of a trigger set. Clients with high reconstruction errors are rejected based on an anomaly detection strategy. While FedCAM achieves very good byzantine robustness compared to concurrent methods in IID scenarios, it fails to achieve good results in the context of non-IID data, as we will see in Section 5. Two main issues cause this failure. First, the CVAE training method proposed in FedCAM has poor generalization in non-IID contexts, as client updates are more heterogeneous. Specifically, clients with the most unique data often fall out of distribution for CVAE and end up with similar error scores to attackers. To overcome this issue, FL-CLEANER employs a different CVAE training process that accounts for the temporal evolution of FL training with warmup epochs and better captures the statistical properties of benign updates using an annealing technique and layer selection. This ensures that all benign client activation maps, even outliers, are correctly reconstructed. Second, FedCAM relies on the mean of reconstruction errors as a dynamic error threshold. This approach can allow the selection of some attackers’ updates for the aggregation step if the average is high enough. It is important to note that retaining even one Byzantine attacker in a round is sufficient to destroy the global model. Moreover, this strategy eliminates around half of the participants in each round. In the non-IID case where clients are less redundant compared to IID scenarios, this impacts the convergence and accuracy of the global model. To address this problem, FL-CLEANER introduces a clustering algorithm that iteratively aggregates and concentrates benign candidates into a trusted cluster, enabling a more refined selection process.
The rest of this paper is organized as follows. Section 2 provides an overview of FL technical background, FL attacks, and useful components of our solution. Section 3 presents the federated context we considered and its threat model. Section 4 introduces our proposal, FL-CLEANER, followed by Section 5, which details our benchmark and provides an analysis of the experimental results. Finally, we conclude the paper in Section 6.
@@ 2.1.1 A Common FL Scenario
In this work, we considered an FL setting where $N$ clients $\{C_{k}\}_{k=1..N}$ with their datasets $\{D_{k}\}_{k=1..N}$ participate in a training session orchestrated by a central server $S$ to train a global model $M_{G}$. During FL round $t$, $S$ scatters the current global model $M_{G}^{t}$ to a subset of $U$ randomly selected clients. Each client updates the received model on its local dataset and sends back the updated model to $S$. The server then uses an aggregation algorithm such as FedAvg [8] to get an updated version of the global model $M_{G}^{t+1}$ from the clients’ updated models $\{M_{u}^{t}\}_{u=1..U}$. The above training steps are then repeated until convergence or a stopping criterion is met.
@@ 2.1.2 Non-IID Data in Federated Learning
Non-IID data hypothesis is considered a challenge for FL defenses for two main reasons. First, a well known problem in FL literature [36, 32] is that it is harder to make the global model converge in this setting, since clients with different data often send contradicting updates. From a defense point of view, this means that it is a risk to misclassify honest clients, requiring precise clients’ udpates filtering. Second, defenses that rely on detecting attackers often look for statistical anomalies to spot these attackers. This is more easy in a IID data context where clients share similar data distribution and thus present similar statistical properties. With non-IID data, clients naturally have different characteristic, and it becomes more difficult to differentiate attackers from natural outliers.
In practice, feature and quality skew are intrinsic properties of data and are hard to simulate experimentally if the dataset does not provide a natural partition. So, most works that address Non-IID FL use a combination of label and quantity skews by choosing a distribution of data over clients that ensures heterogeneity.
2.1.3 Attacks on Federated Learning
Federated Learning is known to be vulnerable to several attacks [32]. These attacks operate at training time and involve an adversary that compromises one or several clients and tries to use them to disrupt the global model. Two main approaches can be distinguished: "model poisoning", where the attacker can directly manipulate the compromised clients’ models and send malicious updates to the server; and "data poisoning", where the attacker indirectly manipulates learning by corrupting the clients’ local datasets. Based on the attacker’s objective, these attacks can further be divided into two sub-classes:
  • Untargeted poisoning attacks - where the attacker’s goal is to drop the global model accuracy on any future inputs by preventing its convergence. It is generally very hard for a machine learning model to recover from a bad convergence, and one attack in an unprotected FL round is often sufficient for an untargeted poisoning attack to succeed [9]. Such attacks are also referred to as "Byzantine attacks".
  • Targeted poisoning attacks - also called "Backdoor attacks", where the attacker aims at preserving the accuracy of the model on its main task while simultaneously training it on a second malicious task, introducing a backdoor. The most common form of backdoor attacks involves inserting a trigger in the training dataset and training the model to output a specific response when the trigger data are given as an input. [37]. Through the control of just some of the FL clients, the attacker may have insufficiently representative data because of the Non-IID context [38], making the backdoor updates easily overwritten by future training rounds, as shown in [38]
@@ 2.2 Conditional Variational Autoencoders (CVAE)
Autoencoders (AEs) [39] are a class of unsupervised generative neural networks that aim to reconstruct a given input from a latent representation encoding. AEs are made of: i) an encoder $f$ that maps input vectors from the initial feature space to a lower-dimensionality representation space called the latent space; and ii) a decoder that retrieves input vectors from their latent representations. The objective of an autoencoder is to minimize the reconstruction error (RE) between the initial and reconstructed vectors. They have been historically been used for dimensionality reduction [40, 39], model pre-training [41, 42], and generative tasks [43]. An important property of AEs is that, when applied to out-of-distribution vectors that deviate from the original training distribution, these deviations are amplified in the latent space, and much higher reconstruction errors are observed. This makes AEs considered as a valuable tool for anomaly detection in a wide range of applications [44].
Autoencoders having a fixed-dimensional latent space show some limitations for datasets with varying complexities. This has motivated the introduction of Variational Autoencoders (VAEs) [45] with a variational latent space based on Bayesian inference methods [46]. Conditional-VAEs (CVAEs) [47] are an extension of VAEs that leverages conditional probabilities to guide the encoding and decoding process following a provided condition. In practice, a CVAE is a VAE with a condition vector $y$ as additional input for both the encoder and decoder (see Fig. 2). CVAEs have been extensively used for a number of applications, including anomaly detection and detecting attackers in federated learning [12, 16, 13].
@@ 2.3 GeoMed and Activation Maps
In a typical deep learning model, data are processed through feed-forward layers. Each hidden layer takes in the previous layer’s output, applies a linear transformation parameterized by a set of weights (e.g., a convolution or weighted sum), and then applies a nonlinearity function. The output of the layer’s weighted transformation is referred to as the activations of that layer. "Activation maps", or the set of activations extracted from a model for a given input, are characteristic of a model and the way it processes data. They have been used for applications such as explainability [49] and model watermarking [50, 51, 52]. It has been observed that in many real-life applications, the activation maps of a model tend to follow an intrinsic distribution [50, 53].